# Towards Single-Lens Controllable Depth-of-Field Imaging via All-in-Focus Aberration Correction and Monocular Depth Estimation

### v1

@@ Appendix C More PSF Estimation Results
Appendix D Monocular Depth Estimation Results
@@
In this section, we present the depth estimation results, where the GT images and aberration images are predicted using the same Monocular Depth Estimation (MDE) model [11]. The qualitative results are shown in Fig. D.1. It is not difficult to find that the depth estimation results of the aberration image are close to that of the GT image. This shows that the MDE model is robust to aberration degradation, so it directly uses the aberration image for MDE to obtain depth information.
Figure D.1: Qualitative comparison of MDE results of the aberration image and the GT image.

### v2

@@ Appendix E Quantitative Analysis of Different state-of-the-art MDE models on the DAMOS Dataset
Appendix F Qualitative Analysis of Monocular Depth Estimation Results for Aberration Images and GT Images
@@
In this section, we present the depth estimation qualitative results, where the GT images and aberration images are predicted using the same metric Monocular Depth Estimation (MDE) model [11].
The qualitative results are shown in Fig. F.1. The depth prediction results for both the aberration images and the GT images exhibit acceptable visual differences, with the depth predictions from the aberration images remaining reliable.
Figure F.1: Qualitative comparison of MDE results of the aberration image and the GT image.
Appendix G Comparison of Aberration Correction Using Different Model Paradigms
We compare the image restoration results of the SR-based method with those of defocus deblurring approaches [50, 51]. As shown in Table G.1, the SR-based method outperforms the defocus deblurring method in terms of metrics. This advantage is rooted in findings from a previous study [16], which demonstrated that the cylindrical network paradigm within the SR-based framework performs better on CAC tasks. Achieving high-quality image restoration results is essential for obtaining single-lens controllable DoF imaging. Additionally, this work highlights the proposed training scheme and depth-aware mechanism, and our framework is not restricted to a specific network paradigm; other paradigms are also applicable.
TABLE G.1: Quantitatively compare the aberration correction capabilities of the defocus deblurring method and our method.
Method Pipeline MOS-S1 MOS-S2
NYUv2 MiddleBury2014 NYUv2 MiddleBury2014
[…]
AIFNet [50] Depth-aware 37.7415 0.9715 0.0373 35.3338 0.9570 0.0763 34.1441 0.9388 0.0745 35.0936 0.9411 0.1055
DRBNet [51] Depth-aware 38.1294 0.9732 0.0359 35.4379 0.9587 0.0759 34.4226 0.9410 0.0735 34.9117 0.9427 0.1059
RSTB [24] Depth-aware 38.8280 0.9758 0.0342 35.8148 0.9616 0.0736 34.8523 0.9450 0.0693 35.3416 0.9446 0.1055
Appendix H Comparison of qualitative results of controllable DoF imaging
To better demonstrate the rendering quality achieved by the proposed framework, we conducted a comparison with an existing rendering-based approach, BokehMe [58]. As shown in Fig. H.1, we first obtain the depth map of an aberration image through the metric MDE model [11] and then restore the image through the CAC module.
Fig. H.1(d) represents the controllable DoF imaging of different lenses generated by our proposed method. As illustrated in Fig. H.1(e), BokehMe’s imaging results achieve different degrees of blur by adjusting the blur parameter $K$, and the blur radius can be calculated by
$r=K\left|d-d_{f}\right|,$ (H.1)
where $d$ denotes disparity (inverse depth) and $d_{f}$ represents the disparity of the focal plane.
To facilitate fair comparison, we chose $K{=}\{5,10,20\}$ respectively to ensure that the blur level is similar to the three lenses represented by Omni-Lens-Field. By closely examining the three images in the far-right column, compared to BokehME, which relies solely on intensity modulation based on blur radius, our method generates blur effects of different lenses by characterizing the PSF of the real lens. This method can accurately reflect the complex effects of optical aberrations (including dispersion).
Figure H.1: Comparison of qualitative results of controllable DoF imaging.
Appendix I Depth distribution of the used datasets
As shown in Fig. I.1, we show the depth distribution of the NYUv2 [69] and Middlebury2014 [70] datasets. Therefore, in this paper, the depth range of the test data is concentrated between $0.7m$ and $10m$ to obtain better comparison results.
Figure I.1: Depth distribution of the NYUv2 and Middlebury2014 datasets.